# Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling

### v8

@@ 2.5.1 Thermodynamic Training Criterion
The effect of tempering is illustrated in Fig. 1, starting with a Gaussian prior that is entirely uninformed of $\bm{x}^{(b)}$. As $t$ increases from 0 to 1, the contribution of the likelihood term in Eq. 29 progressively grows, ultimately leading to the posterior at $t=1$. This final posterior can be interpreted as the prior updated by information from the samples of $\bm{x}^{(b)}$.
@@
$\bm{t}=\left(t_{0},t_{1},\ldots t_{N_{\text{t}}}\right),\quad\quad t_{k}\in[0,1]$ (30)
In Sec. A.4.7, the Steppingstone estimator (SE) introduced by Annis2019 is derived from the discretized Thermodynamic Integral presented by CALDERHEAD20094028, (and provided under Sec. A.4.6). This establishes a connection between the two, which we use to derive the learning gradient and inform our approach to the clustering the temperatures in Eq. 30:
[…]
$\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)=\log(\bm{Z}_{t_{N_{t}}})^{\text{SS}}=\sum_{k=1}^{N_{t}}\log\Bigg(\frac{\bm{Z}_{t_{k}}}{\bm{Z}_{t_{k-1}}}\Bigg)$
[…]
SE and its variants, such as Annealed Importance Sampling (AIS) (neal1998annealedimportancesampling), are commonly used to estimate a model’s evidence without a tractable way to learn the prior, as they lack dependence on prior gradients beyond sampling. However, by recognizing the direct connection between the SE and MLE learning gradients, as proven in Sec. A.4.7, one can simply add the contrastive divergence learning gradient from Eq. 18 to the SE estimator.
This is valid because the SE gradient matches the MLE gradient, and the likelihood gradient is independent of the prior gradient. The posterior expectation of contrastive divergence in Eq. 18 can be estimated with a Monte Carlo estimator using the final samples of ULA.
2.5.2 Unadjusted Langevin Algorithm (ULA) and Replica Exchange
In their implementation of Thermodynamic Integration, CALDERHEAD20094028 estimated the Thermodynamic Integral using the discretization in Sec. A.4.6, using population-based MCMC and Parallel Tempering (PT) (PT; Marinari1992; Hukushima1996) to simultaneously sample from all power posteriors.
Specifically, parallel chains were maintained at each temperature, (or Replica), using Metropolis-Hastings (MH) (13ab5b5e-0237-33fb-a7a8-6f6e4e0d4e0f), to drive local moves for each power posterior. They also proposed a geometric path between the prior and posterior, enabling global swaps between adjacent temperatures. This approach allowed higher temperatures to leverage the efficient mixing of lower temperatures, thereby facilitating a more thorough exploration of the posterior landscape.
We adopt a similar strategy but follow pang2020learninglatentspaceenergybased in using the Unadjusted Langevin Algorithm (ULA) (Rossky1978; Brooks2011) without MH correction. Specifically, we update samples using the forward Euler-Maruyama discretization of the Langevin SDE, as detailed by Roberts2002:
$\bm{z}^{(i+1,\;t_{k})}=\bm{z}^{(i,\;t_{k})}+\eta\nabla_{\bm{z}^{(i,\;t_{k})}}\log\gamma\left(\bm{z}^{(i,\;t_{k})}\right)+\sqrt{2\eta}\bm{\epsilon}^{(i,\;t_{k})},$ (33)
where $\eta$ is the step size, $\bm{\epsilon}^{(i,\;t_{k})}\sim\mathcal{N}(\bm{0},\bm{I})$, and $\gamma_{t_{k}}$ denotes the target power posterior:
$\log\gamma\left(\bm{z}^{(i,\;t_{k})}\right)=\log p_{\bm{f},\bm{\Phi}}(\bm{z}^{(i,\;t_{k})}\mid\bm{x}^{(b)},t_{k})\propto t_{k}\log p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})})+\log p_{\bm{f}}(\bm{z}^{(i,\;t_{k})}).$
ULA is best suited for smooth, unimodal distributions, but annealing with power posteriors can improve sampling from multimodal distributions. Mixing can be further improved by periodically swapping populations between randomly selected adjacent temperatures, subject to:
$r=\frac{p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k+1})}\right)^{t_{k}}p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})}\right)^{t_{k+1}}}{p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k})}\right)^{t_{k}}p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}^{(i,\;t_{k+1})}\right)^{t_{k+1}}},$ (34)
where the proposed swap, $\bm{z}^{(i,\;t_{k+1})}\leftrightarrow\bm{z}^{(i,\;t_{k})}$, is accepted with probability: $\min\left(1,r\right)$.
3 Experiments
Our experiments are designed to highlight the unique advantages of our contributions and compare the proposed strategies for scaling and training KAEM. Radial basis functions (RBFs) (li2024kolmogorovarnoldnetworksradialbasis) were used as the univariate functions in place of the cubic B-spline interpolations proposed by liu2024kankolmogorovarnoldnetworks. Benchmarks are provided in Sec. A.2.
In Sec. 3.1, we show that importance sampling becomes a viable and highly efficient training strategy for simple datasets, due to the low dimensionality of the latent space. We also demonstrate that the model remains effective when strictly adhering to KART with independent energy-based priors, thereby validating the use of KART as a source of structural inductive bias.
In Sec. 3.2, we evaluate the mixture EBM prior, sampled via component-wise ITS and paired with CNN-decoder architectures, which are slightly smaller variants of those used by pang2020learninglatentspaceenergybased. We additionally compare KAEM trained using single-posterior ULA sampling with the MLE criterion against a population-based ULA sampling strategy using the SE criterion. These comparisons assess the practicality of using mixture densities for scaling KAEM and using annealing to improve training for high-dimensional RGB images.
To quantify the fidelity of generated data in a more reliable manner than previous studies in the literature, we used effectively unbiased metrics based on the principles outlined by chong2020effectively. By applying linear regression to extrapolate the values of standard metrics to an infinitely sized sample set, Monte Carlo error was mitigated, (otherwise associated with finite sample sizes).
For Sec. 3.2, we gathered $\overline{\text{FID}}_{\infty}$ and $\overline{\text{KID}}_{\infty}$, which are effectively unbiased estimators of the Fréchet Inception Distance and Kernel Inception Distance, introduced by heusel2018gans; binkowski2021demystifying respectively. The set of sample sizes used for the linear regression were: 1000, 1200, 1400, 1600, 1800, 2000, 2200, 2400, 2600, 2800, 3000.
However, we recognise that quantifying the fidelity of generated data is inherently challenging and not an exact science. While the metrics used in this study have limitations, they align with those commonly found in the literature. Our conclusions are also informed by our subjective perception of the images. Furthermore, no metrics were applied in Secs. 3.1.1 &  3.1.1, as vision metrics rely on the Inception V3 model trained on ImageNet (szegedy2015rethinking), potentially introducing bias against other datasets, especially non-RGB images.
3.1 Importance Sampling
3.1.1 MNIST & FMNIST
The MNIST (mnist) and FMNIST (fmnist) datasets are grayscale datasets that can be modeled with the first likelihood model from Eq. 20. The first collection of benchmarks presented in Sec. A.2 were obtained using the hyperparameters and architecture specified for this experiment in Tab. 4. KAEM was trained for 10 epochs with 50,000 training examples, an importance sample size of $N_{s}=100$, and a batch size of $N_{x}=100$, (corresponding to 5,000 parameter updates). Prior sampling was conducted with ITS and posterior expectations were estimated with IS. The architecture was guided by KART and independent priors were used.
(a) Real samples
(b) No ref. prior
(c) Gaussian ref. prior
(d) Uniform ref. prior
Figure 2: Generated MNIST examples using different reference priors
(a) Real samples
(b) No ref. prior
(c) Gaussian ref. prior
(d) Uniform ref. prior
Figure 3: Generated FMNIST examples using different reference priors
KAEM consistently generated a diverse set of samples when non-restrictive reference priors were used, (Gaussian, uniform, or no reference), thus validating importance sampling, independent energy-based priors, and strict adherence to KART for simple datasets.
Fig. 4 depicts a single latent component of KAEM trained on FMNIST with a Gaussian reference prior, demonstrating KAEM’s potential for latent prior discovery.
Generally, priors were modified minimally when trained with IS, and remained largely aligned with their reference distributions. This suggests that the reference prior used to initialize KAEM guides the structure of the resulting latent space and may provide a useful source of inductive bias or reusable structure.
3.2 Langevin Monte Carlo and Steppingstone estimator
KAEM, when adherent to KART, can be interpreted as a dense or spline-based network. However, for spatial data, convolutional neural networks (CNNs) (LeCun2015Deep), are more suitable as the generator, $\mathcal{G}$. In addition to their computational efficiency through weight sharing, CNNs offer a key advantage over dense networks due to their spatial inductive bias: sequences of convolutional layers implement shift-equivariant transformations that are well-suited for images.
We evaluated KAEM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (37648) and CelebA ($64\times 64$) (liu2015deeplearningfaceattributes). The first likelihood model in Eq. 20 was adopted for both datasets. Due to limited computational resources during this study, (detailed in Sec. A.2), training was shortened to 16,000 parameter updates, (70,000 were used in the study by pang2020learninglatentspaceenergybased), and the CNN decoder architectures used by pang2020learninglatentspaceenergybased were adopted as the generators.
A single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator. Full architectural details are provided under Sec. A.6.1. Posterior sampling was conducted with 40 steps of ULA, and a standard Gaussian reference prior was used, $\pi_{0}=\mathcal{N}(0,1)$. KAEM was trained for 40 epochs with a batch size of $N_{x}=50$ on 20,000 training examples. All remaining hyperparameters are in Tab. 5. Annealing was conducted with $N_{t}=5$ power posteriors and the $p$ scheduling outlined in Sec. A.3.1.
Table 1: Effectively unbiased FID and KID scores (lower is better) for SVHN ($32\times 32$).
MLE SE
$\overline{\text{FID}}_{\infty}=78.41$ $\overline{\text{FID}}_{\infty}=165.79$
$\overline{\text{KID}}_{\infty}=0.0627$ $\overline{\text{KID}}_{\infty}=0.1608$
Table 2: Effectively unbiased FID and KID scores (lower is better) for CelebA ($64\times 64$).
MLE SE
$\overline{\text{FID}}_{\infty}=130.05$ $\overline{\text{FID}}_{\infty}=150.25$
$\overline{\text{KID}}_{\infty}=0.1427$ $\overline{\text{KID}}_{\infty}=0.1637$
(a) Real samples
(b) Maximum Likelihood
(c) Steppingstone Estimator
Figure 5: Generated SVHN ($32\times 32$) examples using mixture prior architecture with inverse transform sampling from the prior and Langevin sampling from the posterior
(a) Real samples
(b) Maximum Likelihood
(c) Steppingstone Estimator
Figure 6: Generated CelebA ($64\times 64$) examples using mixture prior architecture with inverse transform sampling from the prior and Langevin sampling from the posterior
Despite being trained with reduced complexity for a small iteration count, our mixture prior model was able to generate low quality samples of SVHN and CelebA. This demonstrates the potential of component-wise ITS for exact and highly efficient sampling in replacement to the approach of pang2020learninglatentspaceenergybased, which relied on ULA to sample from the prior, thus introducing bias and inefficiency.
Our proposed SE training strategy resulted in worse image metrics for both datasets. However, from our subjective perspective, the generated CelebA images appeared more well-defined, yet less diverse under SE than under MLE. Future work could investigate when SE might be preferred over MLE, as SVHN performed well under the MLE criterion, whereas CelebA subjectively benefited from SE to produce human-perceptible features. For example, a more optimally tuned annealing schedule might have sharpened SVHN features in a manner similar to its effect on CelebA.
Ultimately, we cannot justify the Thermodynamic training strategy due to its substantial computational overhead. With more VRAM than available in this study, it would be possible to parallelize a greater number of temperatures than $N_{t}=5$, which could potentially improve quality, as suggested by the results of CALDERHEAD20094028.
More posterior sampling options, such as modified annealing schedules, a larger number of ULA steps, or the use of autoMALA with bias correction and temperature-adaptive step size tuning may further improve results. Nevertheless, exploring these options might not be worthwhile until the hardware described by exa2024 becomes available, given the added algorithmic complexity, hyperparameter sensitivity, and extended training times involved. A better approach to posterior sampling might involve amortized inference similar to kingma2022autoencodingvariationalbayes.
4 Conclusions
KAEM provides a principled, scalable framework for structured and informed generative modeling. Our results demonstrate that importance sampling can enable more efficient training on simple or low-dimensional datasets. We also show how the Kolmogorov-Arnold representation theorem can inform architectural design and introduce structural bias, allowing KAEM’s architecture to be guided entirely by the size of its latent space. Moreover, we find that the inverse transform method offers advantages over Langevin Monte Carlo sampling for energy-based priors in terms of speed and exact sampling. However, the benefits of annealing and population-based MCMC remain inconclusive, suggesting directions for further work. Future investigations into the practicality of our annealing methods depend on continued hardware advancements, such as those introduced by exa2024.
4.1 Future work
4.1.1 Hardware acceleration
Benchmarks are provided in Sec. A.2, compiled to MLIR/XLA using Reactant (Reactantjl2025) and EnzymeMLIR for automatic differentiation (moses2020enzyme; moses2021enzyme_gpu; moses2022enzyme_parallel).
The univariate, localized nature of B-spline interpolations complicates their parallelization on GPUs, since they involve sparse and recursive computations, so RBF bases were used in replacement. The search operations in Eqs. 1423, and 71 are also difficult to parallelize with GPU kernels, since GPUs are generally ill-suited for search algorithms, as they often rely on conditional logic. Parallelizing conditionals can lead to thread divergence, making GPUs inefficient for these operations. To compile the model with Reactant, these algorithms were instead written using sparse multiplications for conditional masking, which was efficient but suboptimal on GPUs.
Additionally, while ITS enables fast inference and scales when the prior is a mixture model, KAEM’s training with ULA does not scale efficiently on GPUs due to the iterative, gradient-based nature of ULA. GPUs are not well suited to Langevin Monte Carlo methods due to the irregular memory access patterns and bandwidth pressure caused by differentiating with respect to inputs, which requires storing the outputs of each layer within the generator.
Nonetheless, the unique advantages of KAEM justify continued research and development. Moreover, exa2024 introduces the Learnable Function Unit (LFU), a processor designed for efficient evaluation and differentiation of univariate, nonlinear functions in parallel, while offering precise data flow control that can mitigate the challenges of sparsity, recursion, and irregular memory access patterns. The LFU presents a promising avenue for more efficient implementations of ULA, and improved posterior sampling, potentially including bias correction and adaptive step size tuning with autoMALA (bironlattes2024automalalocallyadaptivemetropolisadjusted) to assist when a single step size is insufficient for all power posterior geometries. Even without annealing, replacing ULA with autoMALA may yield a more robust probabilistic model with stronger theoretical guarantees.
KAEM might prove more scalable than conventional generative models, owing to its reliance on finite representations of continuous functions and its compatibility with the LFU.
4.1.2 More opportunities for importance sampling
For SVHN and CelebA, IS proved infeasible as a posterior sampler, so ULA was used for Sec. 3.2. IS is limited by estimator variance, whereas ULA is often limited by mixing and when sampling multimodal distributions, (addressed in this study with annealing).
Therefore, the failure of IS for the more complex datasets might be attributed to prior-posterior mismatch, which rendered the IS estimator ineffective. Moreover, relying on IS restricts the prior to its initial domain, since proposals are generated from the prior using ITS and finite-domain KAN energy functions. In contrast, ULA can explore values of $\bm{z}$ beyond this domain, enabling adaptation of the prior domain within the grid-updating scheme for KANs proposed by liu2024kankolmogorovarnoldnetworks.
A possible remedy for the curse of dimensionality is to learn in a reduced-order space using Principal Component Analysis (PCA) (Jolliffe2016PCA), similar to the approach of kingma2014semisupervisedlearningdeepgenerative, who also applied PCA to train VAEs. To address prior-posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins. Additionally, efforts should be directed toward selecting appropriate initial domains or adapting the domains dynamically.
Additionally, KAEM could be used to learn domain-specific priors that transfer across generator architectures. In this setting, IS may serve as an efficient strategy for rapidly retraining a new generator on a prior that is already aligned with the posterior before retraining.
4.1.3 Expressivity
The requirements for ITS outlined in Sec. 2.1 are quite restrictive, which may limit expressivity at scale. Additionally, it is often desirable to enrich the latent space with more expressive distributions. For example, Normalizing Flows (papamakarios2021normalizingflowsprobabilisticmodeling) achieve this by applying a sequence of reversible transformations that progressively map a simple base distribution into a complex one.
This might be achievable in KAEM by introducing more latent distributions, or revising Eq. 13 into a Mixture of Experts framework (cai2024surveymixtureexperts) where $\alpha_{q,p}$ could be realized as a more flexible kernel to enable different inductive biases. However, this might be expensive for the model by pang2020learninglatentspaceenergybased since its use of Langevin sampling is computationally intense. For example, the previous work of cui2024learninglatentspacehierarchical, which divides their latent EBM space using the diffusion strategy by ho2020denoisingdiffusionprobabilisticmodels, states that sampling is expensive due to their use of MCMC methods.
KAEM could potentially circumvent the additional cost by using IS as a fast sampler, though this would require addressing its current limitations.
Appendix A Appendix
A.1 Codebase
The study was implemented at https://github.com/thezettascale/KAEM.
A.2 Computational benchmarks
These benchmarks were collected for FP32 on a GPU (NVIDIA GeForce RTX 4060) with Julia’s BenchmarkTools package, (BenchmarkTools). Compilation and reverse differentiation were performed using Julia’s Reactant (Reactantjl2025) and Enzyme (moses2020enzyme; moses2021enzyme_gpu; moses2022enzyme_parallel) packages.
Figure 7: Benchmarks for varying latent dimensions, $n_{z}$, while sampling from independent EBM priors with ITS and estimating posterior expectations with IS. The prior had an input size of $n_{z}$ and an output size of $2n_{z}+1$. The generator received inputs of size $2n_{z}+1$, with a hidden size of $4n_{z}+2$ and an output size of $28\times 28$. All other hyperparameters were consistent with Tab. 4 (RBF bases). They reflect the performance of a reverse autodifferentiation of the MLE criterion, Eq. 16.
Figure 8: Benchmarks for varying amounts of temperatures, $N_{t}$, using a mixture of EBM priors (Tab. 6) with ITS sampling. Posterior expectations were estimated with SE and population-based ULA. All other hyperparameters were kept consistent with Tab. 5 (Celeb-A). They reflect the computational cost of performing a single reverse autodifferentiation of the SE criterion, Eq. 31.
A.3 Guidance on tempering
A.3.1 Cyclic annealing
CALDERHEAD20094028 demonstrated that uniformly scheduling the discretized temperatures in Eq. 30 results in unequal contributions of adjacent power posteriors to the Thermodynamic Integral. The success of TI and SE, relies on selecting each $t_{k}$ to ensure sufficient overlap of adjacent power posteriors. This overlap ensures similarity between adjacent power posteriors, as introduced in Sec. A.5.1, which influences the Monte Carlo estimator in Eq. 32, given that SE is an application of importance sampling, (Annis2019).
This overlap is quantified by the KL divergence term in Eq. 50. Maintaining small KL divergences between adjacent power posteriors can be achieved by clustering temperatures in regions where $\mathbb{P}(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi},t_{k})$ changes rapidly. To accomplish this, we follow CALDERHEAD20094028 in adopting a power-law relationship to schedule the temperatures of Eq. 30:
@@ A.3.1 Cyclic annealing
The study by CALDERHEAD20094028 derives an analytic term for an optimal tempering distribution in the context of linear regression models, where the power posterior in Eq. 29 is used to evaluate the plausibility of model parameters in relation to the data, rather than reflecting a latent variable posterior. However, their findings may not hold in the context of KAEM and generative modeling, making $p$ and the temperature schedule difficult variables to tune.
@@ A.3.2 Choosing hyperparameters
$N_{\text{local}}>1$ Determines the number of ULA or autoMALA iterations per power posterior. Although increasing this value improves the quality of samples at each stage, it also raises the computational cost, much like $N_{t}$. It can be kept at a small value, such as $N_{\text{local}}=40$, (based on the recommendations of nijkamp2019learningnonconvergentnonpersistentshortrun; nijkamp2019anatomymcmcbasedmaximumlikelihood regarding short-run Markov chain Monte Carlo).
$N_{\text{unadjusted}}>1$ Controls the number of autoMALA iterations performed without MH adjustments, reflecting the level of exploration at each power posterior. This can be set to 1, as guided by bironlattes2024automalalocallyadaptivemetropolisadjusted.
$p_{\text{init}}>1,$ $p_{\text{end}}\leq 1,$ $N_{\text{cycles}}\geq 0$ Controls the evolution of tempering during training. If the data is expected to have smooth relationships, reduce both $p_{\text{init}}$ and $p_{\text{end}}$. For data with finer details, increase both values to allow for better focus on capturing intricate features. The values used for this study were $p_{\text{init}}=3$, and $p_{\text{end}}=1$, since our experience suggests that power posteriors initially change rapidly closer to the prior. The number of cycles, $N_{\text{cycles}}$, controls the periodicity of cosine scheduling similar to cyclic annealing (fu2019cyclicalannealingschedulesimple).
A.4 Derivations and other identities
A.4.1 Useful proof
@@
$\mathbb{E}_{p_{\bm{\theta}}(\bm{x})}\left[\nabla_{\bm{\theta}}\log p_{\bm{\theta}}(\bm{x})\right]$ $=\mathbb{E}_{p_{\bm{\theta}}(\bm{x})}\left[\frac{\nabla_{\bm{\theta}}p_{\bm{\theta}}(\bm{x})}{p_{\bm{\theta}}(\bm{x})}\right]=\int_{\mathcal{X}}\frac{\nabla_{\bm{\theta}}p_{\bm{\theta}}(\bm{x})}{p_{\bm{\theta}}(\bm{x})}\,\,\,dp_{\bm{\theta}}(\bm{x})$
@@
A.4.2 Marginal likelihood
$\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\left[\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)},\bm{z})\right)\right]$
$=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\left[\log\left(p_{\bm{f}}\left(\bm{z}\right)\cdot p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)\right]$
$=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\left[\log\left(p_{\bm{f}}\left(\bm{z}\right)\right)+\log\left(p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)\right].$ (37)
A.4.3 Maximum likelihood learning gradient

### v9

@@ D.1 Cyclic annealing
The effect of tempering is illustrated in Fig. 9, starting with a Gaussian prior that is entirely uninformed of $\bm{x}$. As $t$ increases from 0 to 1, the contribution of the likelihood term in Eq. 12 progressively grows, ultimately leading to the posterior at $t=1$. This final posterior can be interpreted as the prior updated by information from the samples of $\bm{x}$.
@@
Calderhead and Girolami (2009) demonstrated that uniformly scheduling the discretized temperatures in Eq. 22 results in unequal contributions of adjacent power posteriors to the Thermodynamic Integral. The success of TI and SE, relies on selecting each $t_{k}$ to ensure sufficient overlap of adjacent power posteriors. This overlap ensures similarity between adjacent power posteriors, as introduced in Sec. C.1, which influences the Monte Carlo estimator in Eq. 24, given that SE is an application of importance sampling, (Annis et al., 2019).
This overlap is quantified by the KL divergence term in Eq. 55. Maintaining small KL divergences between adjacent power posteriors can be achieved by clustering temperatures in regions where $p_{f,\Phi}(\bm{z}\mid\bm{x},t_{k})$ changes rapidly. To accomplish this, we follow Calderhead and Girolami (2009) in adopting a power-law relationship to schedule the temperatures of Eq. 22:
@@
The study by Calderhead and Girolami (2009) derives an analytic term for an optimal tempering distribution in the context of linear regression models, where the power posterior in Eq. 12 is used to evaluate the plausibility of model parameters in relation to the data, rather than reflecting a latent variable posterior. However, their findings may not hold in the context of KAEM and generative modeling, making $p$ and the temperature schedule difficult variables to tune.
@@ D.2 Choosing hyperparameters
$N_{\text{local}}>1$ Determines the number of ULA or autoMALA iterations per power posterior. Although increasing this value improves the quality of samples at each stage, it also raises the computational cost, much like $N_{t}$. It can be kept at a small value, such as $N_{\text{local}}=40$, (based on the recommendations of Nijkamp et al. (2019b, a) regarding short-run Markov chain Monte Carlo).
$p_{\text{init}}>1,$ $p_{\text{end}}\leq 1,$ $N_{\text{cycles}}\geq 0$ Controls the evolution of tempering during training. If the data is expected to have smooth relationships, reduce both $p_{\text{init}}$ and $p_{\text{end}}$. For data with finer details, increase both values to allow for better focus on capturing intricate features. The values used for this study were $p_{\text{init}}=3$, and $p_{\text{end}}=1$, since our experience suggests that power posteriors initially change rapidly closer to the prior. The number of cycles, $N_{\text{cycles}}$, controls the periodicity of cosine scheduling similar to cyclic annealing (Fu et al., 2019).
Appendix E Derivations and useful identities
E.0.1 Useful proof
@@
$\mathbb{E}_{p_{\bm{\theta}}(\bm{x})}\left[\nabla_{\bm{\theta}}\log p_{\bm{\theta}}(\bm{x})\right]$ $=\mathbb{E}_{p_{\bm{\theta}}(\bm{x})}\left[\frac{\nabla_{\bm{\theta}}p_{\bm{\theta}}(\bm{x})}{p_{\bm{\theta}}(\bm{x})}\right]=\int_{\mathcal{X}}\frac{\nabla_{\bm{\theta}}p_{\bm{\theta}}(\bm{x})}{p_{\bm{\theta}}(\bm{x})}\,\,\,p_{\bm{\theta}}(\bm{x})d\bm{x}$
@@
E.0.2 Marginal likelihood
$\log\left(p_{f,\Phi}(\bm{x})\right)$ $=\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}\left[\log\left(p_{f,\Phi}(\bm{x},\bm{z})\right)\right]$
$=\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}\left[\log\left(p_{f}\left(\bm{z}\right)\cdot p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\right)\right]$
$=\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}\left[\log\left(p_{f}\left(\bm{z}\right)\right)+\log\left(p_{\Phi}\left(\bm{x}\mid\bm{z}\right)\right)\right].$ (42)
E.0.3 Maximum likelihood learning gradient